OpenStudy (anonymous):

The measurement of the height of 600 students of a college is normally distributed with a mean of 175 centimeters and a standard deviation of 5 centimeters. What percent of students are between 170 cm and 180 cm in height?

OpenStudy (amistre64):

there is an empirical rule that tells us:\[P(-1\sigma\le x\le1\sigma)\approx .6800\] \[P(-2\sigma\le x\le2\sigma)\approx .9500\] \[P(-3\sigma\le x\le3\sigma)\approx .9900\]

OpenStudy (amistre64):

170 < 175 < 180 175-5 < 175 < 175+5 -1\(\sigma\) < 175 < 1\(\sigma\)

OpenStudy (amistre64):

and they want percent, so id change the decimals to percents

OpenStudy (anonymous):

so 68 %?

OpenStudy (amistre64):

id go with 68%, yes

OpenStudy (anonymous):

kk thanks